# Topological BF Theory construction of twisted dihedral quantum double phases from spontaneous symmetry breaking

### v1

@@ Abstract
Nonabelian topological orders host exotic anyons central to quantum computing, yet established realizations rely on case-by-case constructions that are often conceptually involved. In this work, we present a systematic construction of nonabelian dihedral quantum double phases based on a continuous $O(2)$ gauge field. We first formulate a topological $S[O(2)\times O(2)]$ BF theory, and by identifying the Wilson loops and twist operators of this theory with anyons, we show that our topological BF theory reproduces the complete anyon data, and can incorporate all Dijkgraaf–Witten twists. Building on this correspondence, we present a microscopic model with $O(2)$ lattice gauge field coupled to Ising and rotor matter whose Higgsing yields the desired dihedral quantum double phase. A perturbative renormalization group analysis further indicates a direct transition from this phase to a $U(1)$ Coulomb or chiral topological phase at a stable multicritical point with emergent $O(3)$ symmetry. Our proposal offers an alternative route to nonabelian topological order with promising prospects in synthetic gauge field platforms.
@@
Quantum double and Dijkgraaf–Witten twist.— The twisted quantum double $D^{\omega}(G)$ is a systematic construction of certain nonchiral topological orders [17, 8, 18, 19]. The physical data of topological order is completely given by a finite group $G$ and a cohomology class $\omega\in H^{3}(G,U(1))$; the latter encodes the DWT. The physical data refers to the low energy gapped excitations (the anyons) and their statistics: an anyon $a\in\mathcal{A}=\{a,b,c,$...\} is a shorthand for the corresponding excited state $|a\rangle$. The statistics consists of the anyon quantum dimensions $d_{a}$, the fusion coefficients $N^{c}_{ab}$, the $F$-symbol $F^{abc}_{d}$ and the $R$-symbol $R^{ab}_{c}$: the quantum dimensions $\{d_{a}\}_{a\in\mathcal{A}}$ collectively determine the ground state degeneracy (GSD) of topological order on a genus $g$ surface: $\text{GSD}=\sum_{a\in\mathcal{A}}(d_{a})^{g}$, and any symbol $\mathcal{O}^{cd...}_{ab...}$ governs the transition amplitide for incoming anyons ($a,b,$...) and outcoming anyons ($c,d,$...). For $D^{\omega}(G)$, the anyons are labeled by $a=\{(C\ell,\rho)\}$ where $C\ell$ runs over the conjugacy classes of $G$ and $\rho$ the irreducible representations (irreps) of the centralizer of any $g\in C\ell$, see Table 1 for odd $n$ (even $n$ case is left in [19]). The symbols $N^{c}_{ab}$, $F^{abc}_{d}$, and $R^{ab}_{c}$ can all be derived from the algebras of conjugacy classes and irreps of $G$  [^1] The $F$ symbols for the non-twisted quantum double can be chosen to be trivial as this is the usual $D_{n}$ gauge theory in which all the pure gauge fluxes carry bosonic exchange statistics. For this reason, the $F$-symbols and $R$-symbols of $D^{\omega}(G)$ will only depend on the twist, i.e. the 3-cocycle $\omega\in H^{3}(G,U(1))$; and in particular they assume the form of $f^{g_{1}g_{2}g_{3}}_{g_{4}=g_{1}g_{2}g_{3}}$ and $R^{g_{1}g_{2}}_{g_{3}=g_{1}g_{2}}$, depending only on the group elements $g_{i}\in D_{n}$.. The physical data for topological order can be equivalently determined from the modular $S$ and $T$ matrices [21].
@@
$S[O(2)\times O(2)]$ BF theory for dihedral quantum double.— It is known that the $\mathbb{Z}_{n}$ QD can be described by a $U(1)\times U(1)$ BF theory [22, 23, 24]. This construction is based on the fact that $\mathbb{Z}_{n}$ is a subgroup of $U(1)$. Moving to the non-abelian case of dihedral group, since $D_{n}=\mathbb{Z}_{n}\rtimes\mathbb{Z}_{2}$ and $O(2)=U(1)\rtimes\mathbb{Z}_{2}$, it is natural to also expect a $D_{n}$ QD phase to be described by $O(2)$ gauge fields. Concretely, we propose the following gauge theory:
@@
Correspondence between $D^{\omega}(D_{n})$ and $S[O(2)\times O(2)]$ BF theory.— In order to show that $S[O(2)\times O(2)]$ BF theory is a valid description for the dihedral QD, we analyze the states and sectors of the $S[O(2)\times O(2)]$ BF theory on Riemann surfaces: the states are obtained from those of the $U(1)\times U(1)$ BF theory by gauging the symmetry $\mathbb{Z}_{2}^{C}$ defined in (3). Consider a genus $g$ surface $\Sigma_{g}$ with noncontractible loops $\alpha_{i}$, $\beta_{i}$ with i=1,2,,gi=1,2,...,g, where $\alpha_{i}$ and $\beta_{j}$ intersect if and only if $i=j$. On $\Sigma_{g}$, the ground states of a $U(1)\times U(1)$ BF theory are [28, 29, 30, 31]
@@
A one-to-one correspondence between anyons of $D^{\omega}(D_{n})$ and the operators of the $S[O(2)\times O(2)]$ BF theory can also be drawn. On a torus with $g=1$  [^4] To show the correspondence it suffices to restrict the theory to a torus., depending on whether there are twists across of tha gauge fields on the loops $\alpha$ and $\beta$, the total Hilbert space is resolved into four sectors $\mathcal{H}_{\bullet\bullet}$ with $\bullet\bullet=++$, ++-, +-+, and --. The untwisted sector $\mathcal{H}_{++}$ has its states labeled by $\left|p,q\right$$_{++}$, whereas each of the three twisted sectors contains only one state for odd $n$.
@@
for $(p,q)\neq(\pm p^{\prime},\pm q^{\prime})$. Therefore, these operators in $S[O(2)\times O(2)]$ BF theory all have a quantum dimension $d=2$, consistent with that of the $\psi$ anyons in $D^{\omega}(D_{n})$. $W_{p,p}$ are identified with $\psi_{l=p}$ in Table 1, while the operators $\{W_{p,q}\}_{p\neq q}$ are identified with $\{\psi_{l,r}\}$.
@@
where a $T$ matrix calculation gives their topological spins $\theta_{1}=\theta_{\eta}=0$, $\theta_{\sigma_{+}}=\nu/4$, $\theta_{\sigma_{-}}=1/2+\nu/4$, identical to those of the $\eta$ and $\sigma_{\pm}$ anyons in Table. 1. We see that an effect of the $\mathbb{Z}_{2}$ sign $\nu$ is to change the topological spin of $\sigma_{\pm}$ by $1/4$ while preserving that of the other anyons. Physically, the nontrivial $\mathbb{Z}_{2}^{C}$ DWT corresponds to the gluing of a $\mathbb{Z}_{2}^{C}$ SPT state [40, 41, 42, 43], and is reflected in the additional signs of $F$-symbols [5].
@@
The schematic phase diagram for theory (11) is shown in Fig. 2. For sufficiently small Maxwell coupling constant $\kappa$ [Fig. 2 (a)], all gauge fields are deconfined, and the condensation of $s_{\mathbf{i}}$ and $\mathbf{v}_{\mathbf{i}}$ can be separately tuned by $J_{s}/T$ and $J_{v}/T$ respectively, where $T$ is the temperature. This subdivides the phase region with deconfined gauge fields into four phases. Increasing $\kappa$ will confine the discrete gauge fields arising from properly Higgsing the $O(2)$ gauge field, as we analyze below.
Figure 2: Schematic phase diagram of the lattice model Eq. (11) for $n>2$, with schematic RG flows derived from the effective field theory Eq. (12). The relevant RG flow towards the $O(3)$ Wilson-Fisher fixed point is highlighted in orange. The gauge structure of each phase is shown explicitly. The 3D XY and the 3D Ising transitions are marked in red and yellow, respectively. From left to right the three colored planes are phase diagrams for sufficiently small, mediate, and sufficiently large $\kappa$, respectively. The relative position of the $O(3)$ Wilson-Fisher fixed point and the phase diagram for intermediate $\kappa$ is for illustrative purposes only and may differ from the actual situation.
@@
When $J_{s}=+\infty$, the $\mathbb{Z}_{2}$ part of the $O(2)$ gauge field is Higgsed through an Ising transition, and the theory Eq. (11) becomes a lattice Abelian Higgs model with charge-$n$ matter [53, 54, 55, 56]. Increasing $J_{v}/T$ condenses $\mathbf{v}_{\mathbf{i}}$, thereby Higgsing the $U(1)$ gauge structure to $\mathbb{Z}_{n}$; this marks a 3D XY transition from the $U(1)$ Coulomb phase to the $D(\mathbb{Z}_{n})$ phase [54, 55, 56]. Similarly, when $s_{\mathbf{i}}$ is gapped at $J_{s}=0$, tuning $J_{v}/T$ also effects a 3D XY transition between the $O(2)$ Coulomb phase and the $D(D_{n})$ phase. A direct transition between $O(2)$ and $\mathbb{Z}_{n}$ phases can occur when the Ising and the XY transitions meet with each other. The phase diagram is shown in the leftmost panel of Fig. 2.
Stable multicritical point.— It is unnatural to expect a direct transition between $U(1)$ and $D_{n}$ phases, as neither one is the subgroup of the other. However, as we show below through a perturbative renormalization group (RG) analysis, a direct transition can happen at a stable multicritical point where the four stable phases meet, described by the $O(3)$ Wilson–Fisher (WF) universality class with emergent $O(3)$ symmetry.
In the model (11), the $O(2)$ gauge field can be formally integrated out, generating an effective coupling between the vector and the Ising fields. This is legitimate in the deconfined phases, where the Maxiwell coupling $\kappa$ remains weak enough to not invalidate the perturbative expansion in $\kappa$. The effective field theory reads
@@
We first note that, in the absence of the $O(2)$ gauge field, the two matter fields are entirely decoupled in model (11), resulting in a multicritical point where the XY and the Ising transition lines intersect. Since the two transitions are described by $O(2)$ and $O(1)=\mathbb{Z}_{2}$ WF universality classes, the starting point of the perturbative RG analysis for Eq. (12) is the $O(2)\times O(1)$ WF fixed point with $\tilde{\kappa}=0$ and $r_{v,s}>0$. At this fixed point, parameters $t_{v}$, $t_{s}$, and $\tilde{\kappa}$ are relevant, while $r_{v}$ and $r_{s}$ are irrelevant. Among the relevant ones, $t_{v}$ and $t_{s}$ control the XY and the Ising transitions, respectively, whereas $\tilde{\kappa}$ drives the RG trajectory towards the $O(3)$ WF fixed point of Eq. (12), at which $\mathbf{\Phi}_{v}$ and $\phi_{s}$ combine into an $O(3)$ vector and the system exhibits an emergent $O(3)$ symmetry. Since $t_{v}$ and $t_{s}$ perturbations remain relevant in the parameter space of the lattice model, any RG trajectory—as long as it is not the direct trajectory from $O(2)\times O(1)$ to the $O(3)$ WF fixed point—will be driven away from the latter, as illustrated in the left two panels in Fig. 2. This suggests the stability of the multicritical point in the phase diagram. However, if the lattice model allows more relevant perturbations, the multicritical point will be no longer stable, as exemplified in another lattice construction using complex $O(2)$ gauge field given in [19]. We also reiterate that our RG analysis is based on the perturbative expansion in $\kappa$, and might be invalidated with a large enough $\kappa$. Nevertheless, a Monte Carlo simulation (see [19] for details and a detailed RG argument) suggests that for a wide range of $\kappa$ the multicritical point remains stable. Such a direct transition between the $D_{n}$ phase and the $U(1)$ phase at the $O(3)$ WF critical point is similar to the deconfined quantum critical point and SPT transitions [57, 58, 59, 60, 61, 62, 63, 64].
@@
Conclusion and outlook.— In summary, we provide a new route to realizing nonabelian twisted dihedral quantum double phases $D^{\omega}(D_{n})$ through continuous gauge fields: an $S[O(2)\times O(2)]$ BF theory that exactly reproduces the anyon content and modular data of $D^{\omega}(D_{n})$, and a concrete Higgsing construction in an $O(2)$ lattice gauge model with matter fields that realizes these phases microscopically. By tracking Wilson loops and twist operators to anyons and matching their statistics, we establish a one-to-one correspondence that clarifies how discrete dihedral quantum double phases emerge from continuous gauge structure. In the lattice model, we find the dihedral quantum double phase can meet with a $U(1)$ Coulomb phase at a stable multicritical point governed by the $O(3)$ Wilson–Fisher fixed point, enabling a direct continuous transition in analog to deconfined quantum criticality.
Our proposal for realizing nonabelian topological orders is based on continuous gauge theory which distinguishes ours from most existing proposals [8, 14, 15]. Our lattice construction allows us to access a large set of phases and several phase transitions (3D XY, 3D Ising, and $O(3)$ WF) by tuning a small set of parameters. Our construction can be systematized in the realization of more exotic nonabelian topological orders: for example, the point group quantum doubles via Higgsing $O(3)$ or $SU(2)$ gauge fields. Given the success in the experimental creation of synthetic nonabelian continuous gauge fields [47, 67, 68, 69, 70, 71], it is worthwhile to explore similar experiments realizing our proposed field theory (2) and lattice model (11).
Acknowledgments.— ZQG acknowledges Yu-Ping Lin, Hui Yang, Ruihua Fan, Yan-Qi Wang, Salvatore Pace and Xiao-Gang Wen for helpful discussions. CL is grateful for valuable discussions with Jean-Noël Fuchs on related topics. ZQG was funded through the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC02-05-CH11231 (Theory of Materials program KC2301). CL and JEM were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. CL acknowledges support from the EPiQS program of the Gordon and Betty Moore Foundation.

### v2

@@ Abstract
Nonabelian topological orders host exotic anyons central to quantum computing, yet established realizations rely on case-by-case constructions that are often conceptually involved. In this work, we present a systematic construction of nonabelian dihedral quantum double phases based on a continuous $O(2)$ gauge field. We first formulate a topological $S[O(2)\times O(2)]$ BF theory, and by identifying the Wilson loops and twist operators of this theory with anyons, we show that our topological BF theory reproduces the complete anyon data, and can incorporate all Dijkgraaf–Witten twists. Building on this correspondence, we present a microscopic model with $O(2)$ lattice gauge field coupled to Ising and rotor matter whose Higgsing yields the desired dihedral quantum double phase. A perturbative renormalization group analysis further suggests a direct transition from this phase to a $U(1)$ Coulomb or chiral topological phase at a stable multicritical point with emergent $O(3)$ symmetry. Our proposal offers an alternative route to nonabelian topological order with promising prospects in synthetic gauge field platforms.
@@
Quantum double and Dijkgraaf–Witten twist.— The twisted quantum double $D^{\omega}(G)$ is a systematic construction of certain nonchiral topological orders [17, 8, 18, 19]. The physical data of topological order is completely given by a finite group $G$ and a cohomology class $\omega\in H^{3}(G,U(1))$; the latter encodes the DWT. The physical data refers to the low energy gapped excitations (the anyons) and their statistics: an anyon $a\in\mathcal{A}=\{a,b,c,$...\} is a shorthand for the corresponding excited state $|a\rangle$. The statistics consists of the anyon quantum dimensions $d_{a}$, the fusion coefficients $N^{c}_{ab}$, the $F$-symbol $F^{abc}_{d}$ and the $R$-symbol $R^{ab}_{c}$: the quantum dimensions $\{d_{a}\}_{a\in\mathcal{A}}$ collectively determine the ground state degeneracy (GSD) of topological order on a genus $g$ surface: $\text{GSD}=\sum_{a\in\mathcal{A}}(d_{a}/|G|)^{2-2g}$, and any symbol $\mathcal{O}^{cd...}_{ab...}$ governs the transition amplitide for incoming anyons ($a,b,$...) and outcoming anyons ($c,d,$...). For $D^{\omega}(G)$, the anyons are labeled by $a=\{(C\ell,\rho)\}$ where $C\ell$ runs over the conjugacy classes of $G$ and $\rho$ the irreducible representations (irreps) of the centralizer of any $g\in C\ell$, see Table 1 for odd $n$ (even $n$ case is left in [19]). The symbols $N^{c}_{ab}$, $F^{abc}_{d}$, and $R^{ab}_{c}$ can all be derived from the algebras of conjugacy classes and irreps of $G$  [^1] The $F$ symbols for the non-twisted quantum double can be chosen to be trivial as this is the usual $D_{n}$ gauge theory in which all the pure gauge fluxes carry bosonic exchange statistics. For this reason, the $F$-symbols and $R$-symbols of $D^{\omega}(G)$ will only depend on the twist, i.e. the 3-cocycle $\omega\in H^{3}(G,U(1))$; and in particular they assume the form of $f^{g_{1}g_{2}g_{3}}_{g_{4}=g_{1}g_{2}g_{3}}$ and $R^{g_{1}g_{2}}_{g_{3}=g_{1}g_{2}}$, depending only on the group elements $g_{i}\in D_{n}$.. The physical data for topological order can be equivalently determined from the modular $S$ and $T$ matrices [21].
@@
$S[O(2)\times O(2)]$ BF theory for dihedral quantum double.— It is known that the $\mathbb{Z}_{n}$ QD can be described by a $U(1)\times U(1)$ BF theory [22, 5, 23, 24, 25]. This construction is based on the fact that $\mathbb{Z}_{n}$ is a subgroup of $U(1)$. Moving to the non-abelian case of dihedral group, since $D_{n}=\mathbb{Z}_{n}\rtimes\mathbb{Z}_{2}$ and $O(2)=U(1)\rtimes\mathbb{Z}_{2}$, it is natural to also expect a $D_{n}$ QD phase to be described by $O(2)$ gauge fields. Concretely, we propose the following gauge theory:
@@
Correspondence between $D^{\omega}(D_{n})$ and $S[O(2)\times O(2)]$ BF theory.— In order to show that $S[O(2)\times O(2)]$ BF theory is a valid description for the dihedral QD, we analyze the states and sectors of the $S[O(2)\times O(2)]$ BF theory on Riemann surfaces: the states are obtained from those of the $U(1)\times U(1)$ BF theory by gauging the symmetry $\mathbb{Z}_{2}^{C}$ defined in (3). Consider a genus $g$ surface $\Sigma_{g}$ with noncontractible loops $\alpha_{i}$, $\beta_{i}$ with i=1,2,,gi=1,2,...,g, where $\alpha_{i}$ and $\beta_{j}$ intersect if and only if $i=j$. On $\Sigma_{g}$, the ground states of a $U(1)\times U(1)$ BF theory are [29]
@@
A one-to-one correspondence between anyons of $D^{\omega}(D_{n})$ and the operators of the $S[O(2)\times O(2)]$ BF theory can also be drawn. On a torus with $g=1$  [^4] To show the correspondence it suffices to restrict the theory to a torus., depending on whether there are twists across of the gauge fields on the loops $\alpha$ and $\beta$, the total Hilbert space is resolved into four sectors $\mathcal{H}_{\bullet\bullet}$ with $\bullet\bullet=++$, ++-, +-+, and --. The untwisted sector $\mathcal{H}_{++}$ has its states labeled by $\left|p,q\right$$_{++}$, whereas each of the three twisted sectors contains only one state for odd $n$.
@@
for $(p,q)\neq(\pm p^{\prime},\pm q^{\prime})$. Therefore, these operators in $S[O(2)\times O(2)]$ BF theory all have a quantum dimension $d=2$, consistent with that of the $\psi$ anyons in $D^{\omega}(D_{n})$. $W_{p,0}$ are identified with $\psi_{l=p}$ in Table 1, while the operators $\{W_{p,q\neq 0}\}$ are identified with $\{\psi_{l,r}\}$.
@@
where a $T$ matrix calculation gives their topological spins $h_{1}=h_{\eta}=0$, $h_{\sigma_{+}}=\nu/4$, $h_{\sigma_{-}}=1/2+\nu/4$, identical to those of the $\eta$ and $\sigma_{\pm}$ anyons in Table. 1. We see that an effect of the $\mathbb{Z}_{2}$ sign $\nu$ is to change the topological spin of $\sigma_{\pm}$ by $1/4$ while preserving that of the other anyons. Physically, the nontrivial $\mathbb{Z}_{2}^{C}$ DWT corresponds to the gluing of a $\mathbb{Z}_{2}^{C}$ SPT state [38, 39, 40, 41], and is reflected in the additional signs of $F$-symbols [5].
@@
The schematic phase diagram for theory (11) is shown in Fig. 2. For sufficiently small Maxwell coupling constant $\kappa$, all gauge fields are deconfined, and the condensation of $s_{\mathbf{i}}$ and $\mathbf{v}_{\mathbf{i}}$ can be separately tuned by $J_{s}/T$ and $J_{v}/T$ respectively, where $T$ is the temperature. This subdivides the phase region with deconfined gauge fields into four phases. Increasing $\kappa$ will confine the discrete gauge fields arising from properly Higgsing the $O(2)$ gauge field, as we analyze below.
Figure 2: Schematic phase diagram of the lattice model Eq. (11) for $n>2$, with schematic RG flows derived from the effective field theory Eq. (12). The relevant RG flow towards the $O(3)$ symmetric fixed point is highlighted in orange. The gauge structure of each phase is shown explicitly. The 3D $U(1)$ abelian Higgs and the 3D Ising transitions are marked in red and yellow, respectively. From left to right the three colored planes are phase diagrams for sufficiently small, mediate, and sufficiently large $\kappa$, respectively. The relative position of the $O(3)$ symmetric fixed point and the phase diagram for intermediate $\kappa$ is for illustrative purposes only and may differ from the actual situation.
@@
When $J_{s}=+\infty$, the $\mathbb{Z}_{2}$ part of the $O(2)$ gauge field is Higgsed through an Ising transition, and the theory Eq. (11) becomes a lattice Abelian Higgs model with charge-$n$ matter [51, 52, 53, 54, 55]. Increasing $J_{v}/T$ condenses $\mathbf{v}_{\mathbf{i}}$, thereby Higgsing the $U(1)$ gauge structure to $\mathbb{Z}_{n}$; this marks a 3D $U(1)$ transition of an abelian Higgs model [55] from the $U(1)$ Coulomb phase to the $D(\mathbb{Z}_{n})$ phase [52, 53, 54]. Similarly, when $s_{\mathbf{i}}$ is gapped at $J_{s}=0$, tuning $J_{v}/T$ also effects a 3D $U(1)$ transition between the $O(2)$ Coulomb phase and the $D(D_{n})$ phase. A direct transition between $O(2)$ and $\mathbb{Z}_{n}$ phases can occur when the Ising and the $U(1)$ transitions meet with each other. The phase diagram is shown in the leftmost panel of Fig. 2.
Stable multicritical point.— It is unnatural to expect a direct transition between $U(1)$ and $D_{n}$ phases, as neither one is the subgroup of the other. However, a Monte Carlo simulation (see End Matter for details) with monopole suppression suggests that for a wide range of $\kappa$ the multicritical point remains stable. Such a direct transition between the $D_{n}$ phase and the $U(1)$ Coulomb phase is similar to the deconfined quantum critical point and SPT transitions [56, 57, 58, 59, 60, 61, 62, 63]. In what followings we propose a scenario based on a perturbative renormalization group (RG) analysis that a direct transition can happen at a stable multicritical point with an emergent $O(3)$ symmetry where the four stable phases meet.
In the model (11), integrating out the $O(2)$ gauge field generates an effective coupling between the vector and the Ising fields. However, with monopole suppression and in absence of self CS terms, the $O(2)$ gauge field is gapless and the generated effective coupling is long-ranged  [^5] We note that, with DWT, i.e. self CS term in the $O(2)$ gauge field, the Coulomb phases become gapped chiral topological phases, in which integrating out the gauge field generates short-ranged effective coupling. Similarly, without monopole suppression, the continuous gauge fields are also gapped, resulting in short-ranged effective coupling as well. In the RG analysis we adopt a simplified point effective coupling, and we conjecture that it extrapolates to the long-ranged case. The effective field theory reads
@@
We first note that, when $\kappa$ is sufficiently small, the $O(2)$ gauge field is decoupled from the matter fields, and the two matter fields become decoupled in model (11), resulting in a multicritical point where the $U(1)$ and the Ising transition lines, with $O(2)$ and $O(1)=\mathbb{Z}_{2}$ symmetry respectively, intersect with each other. The starting point of the perturbative RG analysis for Eq. (12) is this $O(2)\times O(1)$ symmetric point with $\tilde{\kappa}=0$ and $r_{v,s}>0$, at which parameters $t_{v}$, $t_{s}$, and $\tilde{\kappa}$ are relevant, while $r_{v}$ and $r_{s}$ are irrelevant. Among the relevant ones, $t_{v}$ and $t_{s}$ control the $U(1)$ and the Ising transitions, respectively, whereas $\tilde{\kappa}$ drives the RG trajectory towards a correlated fixed point conjectured to have an emergent $O(3)$ symmetry [65, 66], at which $\mathbf{\Phi}_{v}$ and $\phi_{s}$ combine into an $O(3)$ vector. Since $t_{v}$ and $t_{s}$ perturbations remain relevant in the parameter space of the lattice model, any RG trajectory—as long as it is not the direct trajectory from $O(2)\times O(1)$ to the $O(3)$ symmetric fixed point—will be driven away from the latter, as illustrated in the left two panels in Fig. 2. This supports the stability of the multicritical point in the phase diagram. However, if the lattice model allows more relevant perturbations, the multicritical point will be no longer stable, as exemplified in another lattice construction using complex $O(2)$ gauge field given in [19].
@@
Conclusion and outlook.— In summary, we provide a new route to realizing nonabelian twisted dihedral quantum double phases $D^{\omega}(D_{n})$ through continuous gauge fields: an $S[O(2)\times O(2)]$ BF theory that exactly reproduces the anyon content and modular data of $D^{\omega}(D_{n})$, and a concrete Higgsing construction in an $O(2)$ lattice gauge model with matter fields that realizes these phases microscopically. By tracking Wilson loops and twist operators to anyons and matching their statistics, we establish a one-to-one correspondence that clarifies how discrete dihedral quantum double phases emerge from continuous gauge structure. In the lattice model, we find the dihedral quantum double phase can meet with a $U(1)$ Coulomb phase at a stable multicritical point with emergent $O(3)$ symmetry, enabling a direct continuous transition in analog to deconfined quantum criticality.
Our proposal for realizing nonabelian topological orders is based on continuous gauge theory which distinguishes ours from most existing proposals [8, 14, 15]. Our lattice construction allows us to access a large set of phases and several phase transitions (3D $U(1)$, 3D Ising, and $O(3)$ symmetric) by tuning a small set of parameters. Our construction can be systematized in the realization of more exotic nonabelian topological orders: for example, the point group quantum doubles via Higgsing $O(3)$ or $SU(2)$ gauge fields. Given the success in the experimental creation of synthetic nonabelian continuous gauge fields [45, 69, 70, 71, 72, 73], it is worthwhile to explore similar experiments realizing our proposed field theory (2) and lattice model (11).
Acknowledgments.— ZQG acknowledges Yu-Ping Lin, Hui Yang, Ruihua Fan, Yan-Qi Wang, Salvatore Pace and Xiao-Gang Wen for helpful discussions. CL is grateful for valuable discussions with Nicolas Dupuis, Delamotte Bertrand, Jean-Noël Fuchs, Julien Vidal, Benoît Douçot, and Jean-Bernard Zuber on related topics. We thank Cenke Xu especially for helpful discussions on the lattice model and phase diagram. ZQG was funded through the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC02-05-CH11231 (Theory of Materials program KC2301). CL and JEM were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. CL acknowledges support from the EPiQS program of the Gordon and Betty Moore Foundation.
Appendix A End Matter
Fig. 3 shows the Monte Carlo simulation for the gauge theory (11) on $6\times 6\times 6$ cubic lattice with monopole suppression. Here $|\mathbf{v}|$ and $s$ are the order parameters for the $U(1)$ and $\mathbb{Z}_{2}$ symmetries, respectively. The vertical dashed line in the upper left panel denotes the critical value for the 3D XY transition, $J_{v}/T=0.45422$, and the horizontal dashed line in the upper right panel denotes that for the 3D Ising transition, $J_{s}/T=0.22165$. It can be seen that the multicritical point is always stable in the range of $\kappa$ we consider.
Figure 3: Numerical phase diagram in the $(J_{v}/T,J_{s}/T)$ parameter space for various $\kappa$ from Monte Carlo simulation.